### v1

# From Turbulence to Tranquility: AI-Driven Low-Altitude Network

Abstract
Low Altitude Economy (LAE) networks own transformative potential in urban mobility, emergency response, and aerial logistics. However, these networks face significant challenges in spectrum management, interference mitigation, and real-time coordination across dynamic and resource-constrained environments. After addressing these challenges, this study explores three core elements for enabling intelligent LAE networks as follows machine learning–based spectrum sensing and coexistence, artificial intelligence (AI)-optimized resource allocation and trajectory planning, and testbed-driven validation and standardization. We highlight how federated and reinforcement learning techniques support decentralized, adaptive decision-making under mobility and energy constraints. In addition, we discuss the role of real-world platforms such as AERPAW in bridging the gap between simulation and deployment and enabling iterative system refinement under realistic conditions. This study aims to provide a forward-looking roadmap toward developing efficient and interoperable AI-driven LAE ecosystems.
Index Terms:
Low-altitude economy, unmanned areal vehicle (UAV) communications, machine learning, resource optimization.
I Introduction
Figure 1: Illustration of an AI-driven Low-Altitude Economy ecosystem integrating intelligent communication, sensing, and computing capabilities. Machine learning enhances coordination, autonomy, and efficiency across diverse aerial operations, supported by seamless air-ground collaboration and spectrum-aware infrastructure.
Low Altitude Economy (LAE) network has emerged as one of the most technological trends of the next decade. Defined as using airspace below $30003000$ meters for economic, social, and operational activities, LAE covers various applications: urban air mobility such as air taxis, emergency medical deliveries, precision agriculture, environmental sensing, surveillance, and logistics. Supported by unmanned aerial vehicles (UAVs), electric vertical takeoff and landing (eVTOL) aircrafts, high-altitude balloons, and drone swarms, LAE holds enormous promise to revolutionize transportation systems, reduce urban congestion, improve service delivery, and create new economic opportunities. LAE’s importance can be extended beyond pure innovation by addressing critical global challenges such as climate change mitigation, urban sustainability, and disaster resilience [1]. Therefore, LAE can offer an unprecedented opportunity to transform how cities operate and how services are delivered in both urban and remote areas since it enables faster, cleaner, and more flexible aerial operations.
Despite the promise of LAE, it presents several challenges, particularly in wireless communications. Unlike traditional terrestrial or satellite networks, LAE networks must contend with unprecedented levels of density, mobility, and heterogeneity. Moreover, the airspace is shared by hundreds or thousands of diverse aircraft, often operating under strict latency, reliability, and energy constraints. These factors collectively introduce significant complexity in managing the wireless spectrum, mitigating interference, ensuring safety, maintaining stable connectivity, and coordinating operations.
The integration of these networks with terrestrial systems adds further complexity. To realize the full potential of LAE, seamless coexistence between airborne, terrestrial, and non-terrestrial networks should be provided. Thus, this requires new strategies for dynamic spectrum access, adaptive channel modeling, robust interference management, and collaborative multi-layer network control [2]. Due to complex system architecture, needs, and stringent requirements, traditional rule-based approaches can pose indeadequate to address these challenges. Instead, deep learning (DL) based approaches can play a significant role in meeting the requirements for safe, efficient, and adaptive LAE networks.
DL, particularly distributed learning, offers powerful tools to enable harmonious coexistence of air and ground networks. With federated learning, UAVs can collaboratively sense spectrum usage without central coordination [3]. Reinforcement learning enables aircraft to dynamically adjust their trajectories and resource allocations in response to environmental changes. Generative models such as conditional generative adversarial networks (CGANs) help predict radio conditions and optimize channel utilization [4]. Orchestration of these DL techniques promises to make LAE networks more adaptable, resilient, and scalable.
We introduce a comprehensive architecture designed to integrate communication, computing, and sensing functions in LAE networks, which envisions an intelligent control layer that coordinates air and ground assets using AI-powered models as illustrated in Fig. 1. Combining dynamic spectrum management, edge computing, real-time spectrum sensing, and multi-agent coordination, it aims to create an ecosystem where various platforms can share airspace safely and efficiently. To be able to run such an ecosystem, coordination across layers is essential, particularly in environments experiencing spectrum congestion and high mobility.
In this study, we first investigate challenges in terms of wireless communications and sensing towards fully operational LAE networks and then three fundamental elements for advancing LAE: (i) ML for spectrum sensing and coexistence (ii) ML-assisted resource optimization and trajectory planning, and (iii) the development of standards and real-world validation using our testbeds. For each pillar, we provide a detailed review of open issues, current progress, possible solutions using DL, and future directions. Our ultimate goal is to present a roadmap that bridges the gap between theoretical innovation and practical deployment, enabling the realization of scalable, secure, and intelligent LAE networks.
II Challenges in Low-Altitude Wireless Networks
The growth in LAE systems introduces significant challenges across multiple technical aspects. Although numerous challenges exist, such as air traffic management, energy limitations, security risks, privacy concerns, and regulatory compliance [2], this study focuses specifically on spectrum coexistence, integration of communications, computing, and sensing (C2S), and standardization and testbed, which are addressed in subsections below.
II-A ML in Spectrum Utilization and Coexistence
Due to the highly dynamic and dense nature of low-altitude airspace, optimization of spectrum utilization poses formidable challenges. For example, high mobility among UAVs and eVTOLs leads to Doppler shifts, frequent handoffs, and time-varying multipath fading, which significantly impacts link stability. Another issue is that interference becomes particularly inevitable in urban environments where aerials must coexist with terrestrial base stations, Wi-Fi networks, legacy aviation systems, and radars. Therefore, how to dynamically sense and allocate spectrum resources, predict interference patterns in real time, and balance among competing air and ground users are the main questions remaining open, yet [5].
On the other hand, ML offers opportunities to optimize spectrum utilization and interference management in LAE networks; however, its practical application requires solid solutions for several challenges. One of the most important issues is the non-stationary and spatially correlated nature of the environment. Unlike terrestrial networks, LAE consists of highly mobile platforms such as UAVs and eVTOLs operating at different altitudes and mediums. As a result, it leads to changing signal propagation conditions, including frequent transitions between line-of-sight (LOS) and non-line-of-sight (NLOS) links, and fluctuations in interference levels. ML models trained on localized or static datasets have a risk of failing to generalize across the entire 3D airspace. This might reduce the reliability of spectrum estimation and classification.
Another important challenge is the scarcity of labeled and reliable data required to train robust spectrum sensing models. In LAE environments, acquiring ground truth information about spectrum occupancy or interference is not only costly, but also sometimes impractical. For example, military radar systems or special emergency broadcasts may be difficult to detect or label accurately.
Moreover, cross-layer dynamics create additional complexity. Spectrum utilization in LAE is inherently a cross-layer problem since physical layer measurements such as signal-to-noise ratio (SNR) and Doppler shift interact with medium access control (MAC) and network layers. ML models trained in isolation without timely feedback from other layers may become out of sync with the operational context, which is known as model drift [6]. In practice, synchronization between layers is difficult to achieve because of limited cross-platform communication and the asynchronous nature of distributed LAE nodes. This can reduce spectrum reuse efficiency or harm coordination in dynamic spectrum access.
Finally, the lack of standardized simulation environments and datasets tailored to LAE-specific conditions significantly hinders progress. Most existing ML models for spectrum estimation or interference mitigation are evaluated under simplified or terrestrial scenarios that fail to capture the 3D, multi-agent, and time-varying characteristics of low-altitude networks. In the absence of common metrics, reproducibility is limited, and the generalizability of findings remains uncertain. Therefore, regulation and standardization efforts are essential to establish a reliable ML-based spectrum management infrastructure. Moreover, the opaque decision-making processes of many ML models limit their reliability and regulatory acceptance, particularly in spectrum-sensitive domains such as aviation, where explainability and robustness are prerequisites for deployment. As a result, the integration of ML into LAE spectrum systems is likely to remain an open challenge until these domain-specific constraints are effectively addressed.
II-B Integration of C2S
While effective spectrum utilization and interference management are important to connectivity in LAE networks, they show only one dimension of a broader challenge. The integration of C2S in LAE networks presents a set of technical challenges, such as the need to jointly optimize diverse functions under high mobility and resource constraints. In LAE, UAVs are expected to simultaneously operate as mobile edge computing platforms, base stations, relays, and sensors [7]. This results in competing demands on limited onboard resources such as energy, computational capacity, and, of course, link capacity. Coordinating these functions in real time is particularly challenging owing to changing spectrum conditions and low-latency application demands such as delivery and surveillance.
A critical challenge is the scheduling of C2S tasks across a swarm of heterogeneous aerial nodes. For instance, high-resolution sensing workloads can saturate data pipelines so that communication bottlenecks can delay time-sensitive decisions such as path changes or mission replanning. The lack of unified control systems might cause suboptimal performance and poor resource utilization. Furthermore, heterogeneity in UAV platforms (e.g., varying in computational capacity, sensor types, and antenna systems, etc) makes synchronized data fusion and task assignment further challenging [2].
Finally, a key challenge is the lack of standardized architectures that can support scalable, fault-tolerant C2S integration. Current implementations tend to stack functions vertically rather than holistically designing coordinated workflows with feedback across layers. Bridging this gap requires not only algorithmic developments but also robust system designs that consider the interactions among the physical layer, the network, and sensors.
II-C Standardization and Testbed
Beyond challenges in spectrum management and integrated C2S functionality, the improvement of LAE networks has been significantly limited by the lack of standards and real-world testbeds. Existing standards developed for terrestrial 5G, satellite communications, and manned aviation cannot meet the operational and regulatory demands specific to low-altitude platforms.
In parallel, the limited availability of publicly available testbeds undermines the validation and benchmarking of proposed systems and algorithms. Although simulation remains useful for preliminary analysis, it fails to capture some physical layer nuances of LAE environments, such as altitude-dependent multipath fading, spatiotemporal spectrum utilization, and platform-dependent distortions. As detailed later, to address this, the Aerial Experiments and Research Platform for Advanced Wireless (AERPAW) testbed has emerged as a critical resource enabling controlled experiments with programmable UAVs and software-defined radios (SDRs) in mixed urban-rural landscapes [8]. AERPAW supports the evaluation of AI-driven spectrum access and cross-layer optimization under realistic constraints, bridging theory and practice, and hosts over two dozen datasets based on real-world UAV network environments.
However, several challenges still exist. Reproducing dense urban conditions with high fidelity, simulating interference or jamming scenarios, and scaling the testbed infrastructure to accommodate a variety of aerial platforms, such as eVTOLs or fixed-wing drones, are non-trivial tasks. Moreover, the simulation-to-reality gap persists; ML models trained on digital twins often underperform in real-world deployments due to unpredictable environmental conditions [9]. To accelerate LAE adoption, future efforts should prioritize interoperability across testbeds and the development of standardized evaluation metrics. Additionally, collaboration among researchers is essential for sharing test data, harmonizing spectrum policies, and establishing certification protocols.
III AI-Driven Spectrum Management and Interference Mitigation
Figure 2: Federated learning based spectrum sensing performance results versus (a) transmit power, (b) number of UAVs, and (c) training data per UAV.
Spectrum management and interference detection constitute key enablers for robust LAE network operations. Since aerial systems become denser and more dynamic, the ability to intelligently sense, allocate, and manage spectrum resources is essential. AI techniques, particularly ML, have emerged as powerful tools to address these challenges, offering scalable, adaptive, and data-driven approaches. Accordingly, this section discusses ML-based approaches for spectrum sensing and interference detection.
III-A AI-Enabled Dynamic Spectrum Sensing for Low-Altitude Networks
Dynamic spectrum sensing is a fundamental component for enabling reliable and efficient communication in the highly congested LAE, where aerial and terrestrial users share limited frequency resources. In [3], we proposed an federated learning (FL) method to address the challenges of distributed UAV networks, where centralized solutions are impractical because of power constraints, privacy concerns, and high communication overhead. The proposed system utilizes a lightweight convolutional neural network (CNN) architecture, which is specifically designed to operate on resource-constrained UAV platforms. By processing the real and imaginary components of received signals using one-dimensional convolutions, the system captures temporal signal patterns while trying to maintain computational efficiency. A key contribution of the study is the introduction of the FedSNR aggregation method, extending the conventional federated averaging (FedAvg) approach by incorporating the SNR observed at each UAV into the aggregation process. This approach allows the global model to give greater weight to updates from UAVs experiencing higher-quality signals, thereby enhancing adaptability, which is particularly important in heterogeneous and dynamic environments.
[…]
In summary, the proposed FL framework for UAV-based spectrum sensing offers a privacy-preserving and communication-efficient solution. Also, it significantly improves detection accuracy. By joint utilization of a distributed lightweight CNN model with FedSNR aggregation strategy, the system addresses accurate sensing and resource constraints in LAE networks. These findings provide an important foundation for future research and practical implementations in real-world aerial communication systems, where spectrum-sharing requirements demand adaptive and efficient learning solutions.
III-B ML Techniques for Interference Detection and Mitigation
Beyond spectrum sensing, ML plays an essential role in active interference management. Techniques such as reinforcement learning (RL) can allow UAVs to dynamically adjust their transmission power, channel selection, and flight trajectories to minimize interference and maximize throughput [13]. Also, generative models such as 3D-CGANs [4] can further enhance interference mitigation by predicting spatiotemporal channel occupancy maps, enabling UAVs to proactively avoid congested channels. This predictive capability is valuable in dense urban environments because the radio frequency medium can change rapidly due to UAV mobility and reflections from buildings.
Future research directions can include the development of hybrid models that integrate FL with RL to create systems that combine local adaptability with global coordination. Multi-agent RL techniques enable UAV swarms to cooperatively learn interference patterns [14]. Importantly, practical implementations need to consider the trade-offs between computational complexity, latency, and energy consumption to ensure that ML-based approaches do not overwhelm the constrained resources of aerial platforms.
IV AI-Optimized Resource Allocation and Trajectory Planning
Following intelligent spectrum utilization and interference mitigation methods, efficient resource allocation and trajectory planning are a must for stable LAE. The challenge is that these tasks demand real-time decisions over multidimensional resources while having uncertain and dynamic environments. Unlike terrestrial systems, since LAE networks are characterized by high mobility, strict energy and delay constraints, a holistic design approach jointly considering UAV trajectory control and communication resource scheduling is demanded. Hence, this section discusses the possible solutions regarding joint resource allocation and trajectory planning.
IV-A Reinforcement Learning–Based Resource Management in Dynamic Airspace
Optimizing UAV trajectories is essential for reducing electromagnetic interference and improving power efficiency. Conventional preplanned trajectories often fail in highly dynamic environments where network conditions and interference frequently evolve. Also, optimizing only communication performance without considering power consumption and mobility leads to suboptimal results. There exist essential trade-offs among throughput, delay, and power consumption that should be jointly addressed to obtain effective LAE systems. For example, improving communication throughput by flying airborne closer to users requires additional propulsion energy, while reducing delay by speeding up may also rapidly consume limited onboard energy [14].
Compared to terrestrial base stations, aerial nodes can dynamically reposition to maintain LoS links, adapt to ground user movement, and exploit location-dependent channel conditions. But every repositioning decision also impacts delay and power consumption. For example, hovering closer to users can improve SNR but consumes excessive power for rotary-wing UAVs. Similarly, fixed-wing UAVs that are efficient for wide-area scanning but cannot hover require periodic trajectory planning.
Deep reinforcement learning and trajectory-aware resource optimization might be well-tailored to address these multidimensional trade-offs in practical applications. The RL life cycle for LAE resource optimization starts with the UAV observing its environment, including its position, channel state, and weather. Based on a learned policy, it selects an action like adjusting its trajectory or adjusting transmit power. After executing the action, it receives a reward for performance and stability. This experience updates the policy through iterative learning, often using techniques like experience replay and target networks. Over time, this loop of sensing, acting, and learning enables UAVs to adapt to dynamic airspace conditions efficiently.
For example, in [15], it is shown that the digital twin network (DNT) combining gated recurrent units (GRUs) with value decomposition networks (VDN) enables distributed agents like UAVs and base stations to collaboratively optimize spectrum and computing resources. This approach outperforms independent Q-learning baselines, showing the power of cooperative learning. Another interesting study [13] combines apprenticeship learning and inverse reinforcement learning (IRL) to create optimal policies from expert demonstrations. IRL models capture the intent behind expert actions, which allows UAVs to generalize better than traditional behavioral cloning. This is especially useful in LAE environments where manual design of reward functions is difficult due to complex, dynamic objectives.
IV-B Optimization of UAV Trajectories to Minimize Interference
In multi-UAV networks or dense deployments, the challenge is further amplified due to intra-system or inter-system interference. Optimizing spectrum reuse and user association across systems requires coordinated trajectory planning and power control. RL algorithms, such as deep Q-networks, deep deterministic policy gradients, and multi-agent reinforcement learning, allow autonomous adjustment of the flight paths based on real-time feedback. These algorithms can optimize routes to avoid high-interference zones, align with directional beams, and maintain stable links even in dense urban environments.
Moreover, trajectory optimization can improve fairness and quality of service by balancing coverage across users and reducing interference. Future research should investigate the integration of online learning and transfer learning to enable aerial nodes to generalize trajectory strategies across missions and environments.
V Experimental Validation and Testbed Deployment
Besides the AI-driven frameworks and optimization strategies discussed in the previous sections, real-world validation is essential. Experimental validation not only demonstrates the feasibility of theoretical models but also bridges the gap between simulation and practical deployment and provides critical insights for practical LAE networks.
Figure 3: The AERPAW platform can operate in three experimental modes: physical testbed, sandbox, and digital twin environment.
V-A AERPAW: An Experimental Platform for LAE Research
AERPAW, located in Raleigh, NC, serves as a flagship testbed within the NSF Platforms for Advanced Wireless Research (PAWR) program and is uniquely tailored for LAE experimentation. AERPAW, consisting of programmable UAVs and ground nodes equipped with SDRs, offers a highly flexible platform for controlled wireless experiments in heterogeneous environments [8]. By spanning both urban and rural areas (e.g., Lake Wheeler Field Labs and the NC State Centennial Campus), AERPAW enables reproducible and scalable experiments for spectrum management strategies and trajectory optimization techniques discussed above.
As AERPAW combines live outdoor experimentation and digital twin–based emulation, it provides a powerful framework for researchers to design, simulate, and validate aerial wireless communication scenarios before field deployment. This capability is critical to validate the methods and architectures towards regulation, standardization, and practical usage since it allows for iterative refinement of LAE systems.
As illustrated in Fig. 3, other than simulations that are often used in evaluating LAE networks, the AERPAW platform supports three experimental modes. The digital twin mode enables users to develop and test code virtually on cloud-based nodes, offering flexibility but limited physical realism. The sandbox mode employs real radio hardware in controlled indoor environments, allowing scheduled access to live radios without UAV mobility. The testbed mode executes experiments on outdoor nodes in the field, where AERPAW personnel conduct the tests on behalf of users, without live interaction.
V-B Performance Metrics for LAE Validation
Validating LAE technologies at AERPAW involves a comprehensive set of metrics aligned with the core challenges discussed in Section II. These metrics include but are not limited to signal quality, link availability, outage probability, end-to-end throughput, channel rank for multi-antenna systems, latency, and spectrum occupancy. The testbed’s support for packet-level, IQ-level, and RF-level emulation enables fine-grained analysis of such key performance indicators (KPIs) across a wide range of deployment conditions.
Altitude-dependent measurements of propagation conditions, spectrum occupancy, and other KPIs, especially in sub-6 GHz bands, provide valuable insights into the spatial and temporal dynamics of LAE networks, as well as the suitability of specific spectrum bands for supporting different LAE use cases. Experiments such as those conducted at AERPAW are critical, as traditional simulation tools and theoretical analysis often fail to realistically characterize the KPIs of real-world LAE deployments.
V-C Deployment Insights and Testbed Lessons
Practical experience with AERPAW yields several important lessons for LAE development. First, trajectory-aware communication strategies are indispensable under the changing propagation conditions associated with UAV mobility, altitude, and orientation. Second, bridging the simulation-to-reality gap remains a nontrivial challenge. Models trained in digital twin environments often require calibration with real-world measurements to achieve reliable operation. Finally, the modular design of AERPAW’s portable radio nodes and its abstraction layers enables flexible, future-proof architectures capable of accommodating heterogeneous hardware and regulatory requirements.
Fig. 3 presents representative results that evaluate the occupancy of altitude-dependent spectrum in rural and urban environments for LAE networks. In particular, a helikite is used to collect spectrum occupancy data at various altitudes, locations, and years, between 85 MHz to 6 GHz. The cumulative distribution function (CDF) of average received power for LTE downlink Band 13, grouped by UAV altitude intervals, is provided. At lower altitudes (0–40 m), aggregate signal power measurements in Band 13 are significantly weaker, with most measurements falling below -40 dBm. As altitude increases, especially beyond 80 m, received power improves considerably due to enhanced LOS conditions, with the majority of values exceeding -20 dBm.
Also in Fig. 3, we show heatmaps of spectrum occupancy across multiple frequency bands of relevance to LAE networks, for both urban and rural environments, comparing conditions below and above 50 m altitude. We consider the spectrum to be occupied if the received signal power for a specific altitude and frequency exceeds a set threshold. At lower altitudes, spectrum occupancy exhibits significant variation. In earlier urban measurements, some bands (e.g., ISM) show minimal activity, while others (e.g., Bands 13 and 14) are heavily utilized. At higher altitudes, spectrum usage is consistently denser across almost all bands and years, especially in n41 and ISM bands. These findings highlight the importance of altitude-aware spectrum analysis for effective LAE network planning to support different use cases.
The unique characteristics and statistics of the real-world data, such as for altitude-dependent spectrum occupancy, underscore the importance of AI-driven solutions presented in this study. Such data-driven AI solutions can be used to improve spectrum sensing, interference mitigation, resource allocation, and trajectory optimization for LAE networks. They also highlight the need for integrated C2S frameworks and standardized testbed practices. AERPAW plays a critical role in data generation and standardization efforts for LAE networks. The testbed can serve as a validation platform for emerging protocols related to spectrum coexistence, trajectory planning, and resource optimization. By supporting standardized interfaces and benchmarks, AERPAW can help ensure that AI-enhanced solutions are not only technically feasible but also interoperable and aligned with evolving global standards.
VI Open Issues
Moving forward, several open research directions and technical challenges must be addressed to fully realize the potential of AI-driven LAE networks. The primary concern is the integration of explainable AI or trustworthiness techniques into decision-making frameworks. As AI models impact mission-critical operations, it becomes important to ensure transparency and interpretability in their decisions, especially under uncertain or adversarial conditions.
Another issue is the need for developing unified metrics for performance evaluation of AI models in LAE scenarios. Existing datasets and performance metrics can fail to capture the unique dynamics of low-altitude airspace, such as 3D mobility, heterogeneous systems, and dynamic spectrum availability. Realistic simulation environments and standardized evaluation protocols can produce meaningful progress and fair comparisons between different approaches.
Furthermore, establishing interoperable standards between air, space, and terrestrial systems remains a challenge. Cooperation with satellite systems has been studied in the literature, but more efforts are needed for seamless integration that requires well-defined protocols and spectrum policies.
Last, practical concerns such as energy-efficient AI models, robust operation in harsh environments, and real-time adaptation to dynamic conditions require further investigation. Experimental testbeds, FL frameworks, shared datasets, and open-source platforms can help bridge the gap between theoretical advances and real-world deployment. These efforts would also improve community-driven innovation and reduce the risks of commercial adoption of LAE technologies.
VII Concluding Remarks
The emergence of LAE networks presents an exciting frontier in wireless communications as offering transformative capabilities across several domains such as aerial transportation, environmental monitoring, and logistics. However, realizing these promising capabilities requires overcoming challenges in communication, sensing, and system coordination within highly dynamic and resource-constrained airspace environments. In this study, we provide a comprehensive exploration of the key subsystems required to support the vision of AI-driven LAE networks. Along with detailing the challenges towards full operable LAE networks, we identify and discuss three critical issues as follows: spectrum sensing and coexistence strategies empowered by ML, RL-driven resource allocation and trajectory planning, and the standardization and experimental validation of proposed solutions through real-world testbeds such as AERPAW.
We indicate that ML-based approaches can provide significant promise in adapting to the non-stationary, heterogeneous conditions of LAE. For instance, FL models, particularly those incorporating SNR-aware aggregation schemes, can support decentralized and privacy-preserving spectrum sensing while maintaining high accuracy in evolving environments.
We also highlight the role of experimental platforms like AERPAW in narrowing the gap between theoretical innovation and practical deployment. Through digital twins, sandbox environments, and live testbeds, AERPAW enables validation of AI-supported LAE solutions under realistic conditions. These efforts are essential not only for technical development but also for ensuring regulatory compliance, interoperability, and robustness across diverse deployment scenarios.
In summary, AI-driven LAE networks show a convergence of wireless communications, edge intelligence, and autonomous control. By developing ML-based methods, validating them through realistic experiments, and aligning them with standardization efforts, the wireless research community can shape the next generation of safe and intelligent aerial systems.
Biographies
KÜRŞ̧AT TEKBIYIK [StM’19, M’25] (ktekbiyik@gmail.com) received his Ph.D. degree in telecommunications engineering from Istanbul Technical University, Istanbul, Türkiye. His research interests include machine learning applications in wireless communications, terahertz communications, and non-terrestrial networks.
AMIR HOSSEIN FAHIM RAOUF [StM’16] (afahimr@ncsu.edu) is currently pursuing a Ph.D. degree in Electrical Engineering at North Carolina State University, Raleigh, NC, USA. His main research areas are wireless communications, visible light communications, and quantum key distribution.
İSMAİL GÜVENÇ [F’20] (iguvenc@ncsu.edu) received his Ph.D. in electrical engineering from the University of South Florida in 2006. He is currently a professor at NC State University, Raleigh, NC.
MINGZHE CHEN [StM’15, M’19, SM’24] (mingzhe.chen@miami.edu) is currently an Assistant Professor with the Department of Electrical and Computer Engineering at University of Miami. His research interests include federated learning, reinforcement learning, virtual reality, unmanned aerial vehicles, and Internet of Things.
GÜNEŞ̧ KARABULUT KURT [StM’00, M’06, SM’15] (gunes.kurt@polymtl.ca) is Professor of Electrical Engineering at Polytechnique Montréal, Montréal, QC, Canada. She is a Marie Curie Fellow and has received the Turkish Academy of Sciences Outstanding Young Scientist (TÜBA-GEBIP) Award in 2019. She received her Ph.D. degree in electrical engineering from the University of Ottawa, ON, Canada.
ANTOINE LESAGE-LANDRY [StM’16, M’20, SM’24] is an Associate Professor in the Department of Electrical Engineering at Polytechnique Montréal, QC, Canada. His research focuses on decision-making in complex engineering systems like electric power systems and wireless communication networks.